OpenStudy (anonymous):

Charles cuts out a semicircle with a radius of 5 inches from a piece of paper. He then curls it into a cone by joining the two radii on the edge of the semicircle together. What is the lateral surface area of the resulting cone? Round your answer to the nearest hundredth.

OpenStudy (kropot72):

What is the formula for the area of a circle?

OpenStudy (anonymous):

 Pir^2

OpenStudy (anonymous):

It is the area of the semi circle. \[\frac { \pi r^2} 2=\frac { 25 \pi} 2=39.27 \]